OpenStudy (anonymous):

how do i solve using substitution 5x+2y=-17 -4x+y=37

OpenStudy (anonymous):

http://www.twiddla.com/536745

OpenStudy (anonymous):

go there to see work

OpenStudy (gw2011):

-4x+y=37 y=4x+37 Substitute into 5x+2y=-17 5x+2(4x+37)=-17 5x+8x+74=-17 13x=-91 x=-7 Subsitiute x=-7 into -4x+y=37 (-4)(-7)+y=37 28+y=37 y=9 Therefore, x=-7 and y=9 You can substitute x=-7 and y=9 into each of the equations to verify the correctness of the answer